A and B are independent events. P(A)= 0.3, and P(A or B)= 0.8. Find p(B).
Answer by jim_thompson5910(35256)

"A and B are independent events" which means
P(A and B) = P(A)*P(B) = 0.3*x = 0.3x

For now, P(B) is unknown. We'll just leave it as x for now.

P(A or B) = P(A) + P(B) - P(A and B)
0.8 = 0.3 + x - 0.3x

Solve the equation above for x to get your final answer. I'll let you take over and do this part.